There are 8 students.the teacher wants to divide them into groups of two.how many combinations are there?
Found 2 solutions by robertb, sudhanshu_kmr:
Answer by robertb(5830)

8C2*6C2*4C2*2C2 = 2,520.

Answer by sudhanshu_kmr(1152)


here order of group is not to be considered...


so, no. of ways = 8C2*6C2*4C2*2C2/ 4! = 2520/24 = 105